# Carbon-based light emitting defects in different layers of hexagonal boron nitride

### v1

@@ Abstract
Substitutional carbon defects in hexagonal boron nitride (hBN) have garnered significant interest as single photon emitters (SPEs) due to their remarkable optical and quantum properties. An intriguing property of these defects is that they can be spin-active (), even if weakly interacting. Employing density functional theory (DFT) calculations, we demonstrate that two monomers of C-based defects of the same species can exhibit a stable triplet spin state at room temperature, even when they are separated nm, if they reside in different layers. The zero-phonon line (ZPL) energy of C defects in different layers lies within eV range. Also, we found defects that deviate from the typical phonon replica patterns, potentially explaining the observed phonon replicas in yellow emitters in hBN.
@@ I Introduction
This article aims to investigate SPEs when forming a C dimer but residing in different hBN layers. Such a possibility has not been addressed before, even though defects comprising an interstitial C atom have been considered [43]. The paper is structured as follows: Section II details the methods employed for the calculations, Section III presents our main results, including the SPE’s spin state - likely to be a triplet - and their photoluminescence spectra. Finally, Section IV explains the large energy difference between triplet and singlet states for some dimers, the phonon sidebands shape, and we relate our findings to experimental results on spin-active SPEs.
@@ II.1 Computational Methods
We used density functional theory (DFT) as implemented in the Vienna ab initio simulation package (VASP) [44, 45, 46, 47]. A hexagonal supercell was employed together with a single k-point (). We used an hBN bilayer to simulate defects in adjacent layers. If the defects are in non-adjacent layers, we used a hBN trilayer. The connection with the bulk is almost immediate; the additional layers should have (i) a lesser effect on the ZPL energy due to enhanced screening [48, 49, 50] and a shift in the ZPL due to electrostatic potential [51]. We used the AA’ stacking (B, N atoms on top of N, B), as found in most hBN samples [52].
@@
We used the SCAN exchange-correlation functional [56] to relax the hBN structure and to calculate phonons. Although this functional can provide an accurate description of the geometry, it could fail to accurately describe the defect levels [57]. Once the geometry was obtained with SCAN, we used the Heyd-Scuseria-Ernzerhof (HSE06) functional to obtain the energies and the optical transitions at that geometry [58, 59]. This procedure was used for the ground and excited states. To study the excited state of each configuration, the SCF method was used [60, 61]. This method has a problem of correctly describing the excited state, and there are some corrections to improve the ZPL value [7, 62]. The most interesting systems in our study have a triplet ground state and involve transitions with the valence or conduction bands, so getting the correct Slater determinant of the excited state is more involved. In the Supplementary Material, we bound this error in the ZPL energy to eV or lower. In all cases, we restricted our study to neutral defects [32].
@@
where is the frequency of the phonon -th vibrational mode and is the Planck Constant. From here, the optical spectral function and the phonon sideband can be directly determined [66]. The only free parameter is a thermal-like broadening of , which we set to 25 meV (room temperature).
@@ III Results
All of the defects studied are substitutional C-dimers, with both C atoms in different layers. The cases of CNCB, CNCN and CBCB were studied. In all dimers, the C atom in the bottom layer was kept fixed while we changed the position of the C atom in the top layer, see Fig. 1A. For clarity, we labeled the in-plane position () of one C atom with respect to the other with the letters ‘a’-‘g’. In the configuration ‘a’, both C atoms are on top of each other. The in-plane projection of the C-C distance is in Table 1. The C atoms can belong to adjacent layers () or can be separated by an inner layer (); see Fig. 1B-C. Thus, a given dimer defect can be identified by CXCY–, with B, N.
![Refer to caption](x1.png)
Label | Distance (Å) | Distance C-C (Å) | |
a | |||
b | |||
c | |||
e | |||
f | - | ||
g | - |
Almost all configurations with C defects of the same species, i.e. CNCN and CBCB, have a triplet ground state () regardless of their relative position or whether they belong to adjacent layers or not. As shown in Fig. 2, the triplet is approximately 0.5 eV more stable than the singlet state; these defects will remain spin-active well above room temperature. The only exception is CNCN–b. In this particular defect, both atoms attract each other, moving out of the plane, and they form a covalent bond resembling a sp3 hybridization.
On the other hand, for defects of the type CNCB, the ground state is always a singlet. In Section IV, we discuss why the singlet and triplet spin states are favored for different defect types.
![Refer to caption](x2.png)
@@
CBCB–b | 8.65 | 3.45 |
CBCB–e | 8.66 | 3.46 |
CBCB–a | 8.60 | 3.40 |
CBCB–c | 8.73 | 3.53 |
CBCB–f | 8.59 | 3.39 |
CBCB–g | 8.75 | 3.55 |
CNCN–b | 2.80 | 8.00 |
CNCN–e | 3.67 | 8.87 |
CNCN–a | 3.65 | 8.85 |
CNCN–c | 3.63 | 8.83 |
CNCN–f | 3.66 | 8.86 |
CNCN–g | 3.65 | 8.85 |
CNCB–a | 4.50 | 4.50 |
CNCB–c | 4.47 | 4.47 |
CNCB–b | 4.90 | 4.90 |
CNCB–e | 4.94 | 4.94 |
The formation energy of the CXCY defects is provided in Table 2. Regardless of whether the defects reside in the same or different layers [31], the formation energy is primarily determined by the environment (i.e. the chemical potential). In the most favorable environment (N-rich for CBCB, N-poor for CNCN), eV. Conversely, increases significantly to 8.5-9.0 eV in the least favorable environment. Also, details such as the C–C distance have a minimal effect on the formation energy. The only exception is CNCN–b where a covalent C–C bond develops. Note that this defect has the smallest formation energy among all CNCN defects, which is coherent with a partial covalent bond between both C atoms.
Instead, for CBCN, the formation energies are over eV. This is substantially larger than the dimer C2 (i.e CN, CB adjacent, in the same layer), with eV [32]. As a comparison, topological CNCB defects (i.e. Stone-Wales-like) have a formation energy even higher than the CNCB defects studied here [67], but in these defects decreases quickly with strain. Indeed, biaxial strain can substantially lower the relative formation energy of CB or CN depending on whether it is compressive or tensile [68]. While a large strain is not expected to be present in the bulk, C defects can act as stress valves near grain boundaries [67]. The thermal stability of C defects is determined not only by but also by the energy barriers preventing them from migrating. Babar et al. [67] calculated the barrier energy separating a Stone-Wales C2 dimer from the standard C2 dimer; it surpasses 6 eV, as strong covalent bonds need to be broken. In the case of the CXCY–, the same applies, and we do not expect any migration induced by temperature.
![Refer to caption](x3.png)
The ZPL energy for defects with triplet ground state, CNCN and CBCB, is shown in Fig. 3. When an intermediate layer of hBN separates C atoms, the value of ZPL seems to quickly converge to eV for CNCN and eV for CBCB. When the C atoms are on top of each other or very close, the value of the ZPL decreases. As CNCN–b has a C–C bond, breaking would require a photon energy much larger than the hBN band gap; hence we didn’t calculate its optical excitation. The donor-acceptor-like defects, CNCB– have a ZPL eV, which almost coincides with the ZPL of non-adjacent CNCB when residing in the same layer [25].
![Refer to caption](x4.png)
@@
Two types of photoluminescence spectra were found, see Fig. 4A. The one associated with the defect CBCB–c represents most of the defects studied here; there are high-energy phonon replicas as typically found in C-based defects in hBN. However, there is one exception, CBCB–a, with prominent phonon replicas at 62 meV. In the next section, we will discuss the origin of both types of phonon sidebands found.
@@ IV.1 Triplet or singlet spin ground state
Almost regardless of distances or whether defects belong to adjacent layers or are separated by an additional hBN layer, defects of the same type (i.e. CBCB, CNCN) have a triplet ground state, well separated in energy from the singlet state ( eV). The same types of defects can have a triplet ground state when embedded in the same layer, but in that case, the singlet and triplet are nearly degenerate [31]. Instead, when the defects belong to a different type (CBCN), the ground state is a singlet.
The relation between the spin state and the type of defect can be understood by looking at the energy of the defect levels within the fundamental gap of hBN. CB has donor-like defect states (close to the conduction band). In contrast, CN is acceptor-like (defect states closer to the valence band). This is illustrated in panels A and D of Fig. 5, where the configurations CBCB–a and CBCN–b are used as examples.
Panel A of Fig. 5 shows the occupied and empty defect levels of a CNCB defect; they are located around CN and CB, respectively (Fig. 5B, C). This directly affects the total spin of the configurations with different types of defects, as it is energetically more favorable for the two extra electrons of the carbon to form a singlet and localize around the CN defect.
On the other hand, defects of the same species - such as CBCB - are degenerate, and when close enough, their wavefunctions are linear combinations of both defect states, as shown in Fig. 5E-F. Also, there is a non-negligible overlap contribution on the B and N atoms along the line joining both C atoms. As the CB states are degenerate, there is an important overlap of the wavefunctions of both occupied levels. This leads to an exchange energy, lowering the energy of the configuration when the total spin is a triplet. In contrast, when the total spin is a singlet, the exchange energy is null, making the energy of the singlet configuration higher than that of the triplet configuration.
![Refer to caption](x5.png)

### v2

# Optical Properties and Spin States of Inter-layer Carbon Defect Pairs in Hexagonal Boron Nitride: A First-Principles Study

@@ Abstract
Substitutional carbon defects in hexagonal boron nitride (hBN) are prominent single photon emitters (SPEs), and their potential for spin activity () is particularly intriguing. While studies have largely focused on intra-layer defects, we employ density functional theory (DFT) to investigate inter-layer dimers of identical carbon species (CXCX). We demonstrate that these CXCX pairs can exhibit a stable triplet spin state at room temperature when closely spaced (e.g., within 3.5-7.1 Å) across hBN layers. As their separation increases beyond this range (e.g., Å), they transition into weakly interacting pairs, characterized by singlet-triplet degeneracy. This regime is predicted to result in a very small zero-field splitting for the triplet manifold, offering a potential explanation for certain optically detected magnetic resonance (ODMR) signals. The zero-phonon line (ZPL) energy of these inter-layer CXCX pairs is found to be practically monochromatic and within the visible range. Furthermore, we identify specific CBCB inter-layer configurations exhibiting atypical low-energy phonon replicas due to out-of-plane vibrational coupling, a finding that may clarify the vibronic structure of other hBN emitters, such as the ‘yellow emitters’.
@@ I Introduction
This article aims to investigate SPEs when forming a C dimer but residing in different hBN layers. Such a possibility has not been addressed before, even though defects comprising an interstitial C atom have been considered [43]. The paper is structured as follows: Section II details the methods employed for the calculations, Section III presents our main results, including the SPE’s spin state, likely to be a triplet, and their photoluminescence spectra. Finally, Section IV explains the large energy difference between triplet and singlet states for some dimers, the phonon sidebands shape, and we relate our findings to experimental results on spin-active SPEs.
@@ II.1 Computational Methods
Density functional theory (DFT) calculations were performed using the Vienna ab initio simulation package (VASP) [44, 45, 46, 47]. We utilized a hexagonal supercell of hBN with -point sampling. Defects in adjacent layers were modeled using an hBN bilayer, while an hBN trilayer was employed for defects in non-adjacent layers. These multi-layer models aim to approximate bulk hBN; additional layers are expected to influence the zero-phonon line (ZPL) energy through (i) enhanced dielectric screening [48, 49, 50] and (ii) modified electrostatic potentials [51], with these influences diminishing as layer thickness increases towards the bulk limit. The AA’ stacking (B and N atoms on top of N and B, respectively), characteristic of most hBN samples [52], was adopted.
@@
The SCAN exchange-correlation (XC) functional [56] was employed for structural relaxation, phonon calculations, and optical transition energies. The suitability of SCAN for defect levels is supported when the host’s band gap is reasonably described [57], as is the case here (SCAN-predicted hBN gap: 5.0 eV; experimental value eV). While the Heyd-Scuseria-Ernzerhof (HSE06) functional[58] is a common alternative, its results are highly sensitive to dielectric screening, which is significantly reduced and anisotropic in 2D materials like hBN. This sensitivity is pronounced for donor-acceptor defects with strong electrostatic interactions. Consequently, applying HSE06 to our bilayer and trilayer models, which inherently possess different screening characteristics, yields results whose direct comparison or extrapolation to bulk behavior is complex. Nevertheless, HSE06 optical transition energies, calculated using the SCAN-relaxed geometries (a common practice where less demanding functionals provide reliable structures [59]), are reported in the Supplementary Material. On the other hand, the functional of Perdew-Burke-Ernzerhof [60] failed to describe the ground state of the defects at large distances, providing a spurious triplet state. Even worse, when using the PBE’s wavefunctions as seeds for more expensive calculations, the error is propagated.
To study the excited state of each configuration, the SCF method was used [61, 62]. This method has a problem of correctly describing the excited state (i.e. mixed spin states are employed instead of well-defined spins), and there are some corrections to improve the ZPL value [7, 63]. The most interesting systems in our study have a triplet ground state and involve transitions with the valence or conduction bands, so getting the correct Slater determinant of the excited state is too involved to be practical. In the Supplementary Material, we bound this error in the ZPL energy to eV or lower. In all cases, we restricted our study to neutral defects [32]. Calculation of the single-photon emission rates would involve the computation of electric dipole transition moments between the involved excited and ground electronic states, which is beyond the scope of the current work.
@@
where is the frequency of the phonon -th vibrational mode and is the Planck constant. The total Huang-Rhys factor is . The photoluminescence spectral lineshape was then constructed using the generating function formalism, which builds the vibronic progression based on , the set of , and [67]. Finally, the calculated stick spectrum (ZPL and phonon replicas) was convoluted with a Gaussian function with a Full Width at Half Maximum (FWHM) of 25 meV. This phenomenological broadening accounts for homogeneous lifetime effects, inhomogeneous broadening, and approximates thermal effects at room temperature, allowing for comparison with experimental spectra.
From here, the optical spectral function and the phonon sideband can be directly determined [67]. The only free parameter is a thermal-like broadening of , which we set to 25 meV (room temperature).
@@ III Results
All of the defects studied are substitutional C-dimers, with both C atoms in different layers. The cases of CNCB, CNCN and CBCB were studied. In all dimers, the C atom in the bottom layer was kept fixed while we changed the position of the C atom in the top layer, see Fig. 1A. For clarity, we labeled the in-plane position () of one C atom with respect to the other by the in-plane projection of their distance, taking the nearest neighbor distance as ‘1’ (i.e. is equivalent to 1.45 Å), as found in a similar context for labeling C-defects [25]. To convert to times the lattice parameter, it is needed to multiply it by . The conversion of to Å is given in Table 1. The C atoms can belong to adjacent layers () or can be separated by an inner layer (); see Fig. 1B-C. Thus, a given dimer defect can be identified by CXCY–, with B, N.
| Distance (nearest neighbors) | Distance (Å) | Distance C-C (Å) | |
Defects of the type CNCB are consistently found in our calculations to exhibit a singlet ground state (see Fig. 2). This observation aligns with established expectations: CN sites, characterized by defect energy levels close to the valence band maximum, act as acceptors. Conversely, CB sites are understood to behave as donors. Consequently, a charge transfer from CB to CN is anticipated, irrespective of their separation, leading to effectively charged C and C centers. In the dissociation limit, as direct quantum mechanical interactions (such as exchange coupling) between the two carbon atoms vanish, their distinct electronic natures –CB as a donor and CN as an acceptor– remain the dominant factor. Overall, the system will naturally remain a singlet, consistent with our findings.
In contrast, carbon defects of the same species, i.e., CNCN and CBCB, generally exhibit a stable triplet ground state when the defects are relatively close. This tendency is typically observed for separations up to (an in-plane distance of Å) and (a vertical separation of two layers, resulting in a total distance of approximately 7Å), as detailed in Fig. 2. However, for CBCB defects in adjacent layers (), the spin state at shorter distances can be more sensitive to the specific geometric configuration.
The spin state of the defect pair is intrinsically linked to the symmetry of the two-electron spatial wavefunction: a singlet state is associated with a spatially symmetric wavefunction, while a triplet state corresponds to a spatially anti-symmetric wavefunction. This spatial symmetry dictates the charge distribution of the pair; for instance, an anti-symmetric wavefunction (triplet) possesses a nodal plane or region of zero electron density at the midpoint between the defects, whereas a symmetric wavefunction (singlet) typically has finite electron density there. This difference in charge distribution influences the electrostatic interaction with the polar hBN lattice, a mechanism previously invoked to explain the pronounced stability of the triplet state for certain CNCN configurations within the same hBN layer (h=0) [31].
As the defects are moved further apart, the significance of the wavefunction’s symmetry at the midpoint diminishes because the overall wavefunction amplitude in this central region approaches zero. In this dissociation limit, the C atoms evolve into weakly interacting S=1/2 paramagnetic centers [42]. Consequently, the exchange interaction between them vanishes, leading to the singlet and triplet states becoming effectively degenerate, as anticipated for two distant, uncorrelated S=1/2 spins.
A notable exception to the general triplet preference for nearby CNCN defects is the CNCN–(1,1) configuration. In this case, the two CN atoms experience a strong mutual attraction, causing them to displace significantly from the hBN plane and form a direct covalent C-C bond. This bonding scenario, which locally resembles sp3 hybridization, results in a singlet ground state, i.e. consistent with the spin-pairing typically observed upon formation of a strong covalent bond.
@@
| CBCB– | 9.05 | 3.85 |
| CBCB– | 9.08 | 3.88 |
| CBCB– | 9.08 | 3.88 |
| CBCB– | 9.08 | 3.88 |
| CBCB– | 9.07 | 3.87 |
| CBCB– | 9.07 | 3.87 |
| CBCB– | 9.06 | 3.86 |
| CBCB– | 9.06 | 3.86 |
| CBCB– | 9.05 | 3.85 |
| CBCB– | 9.06 | 3.86 |
| CNCN– | 2.88 | 8.08 |
| CNCN– | 3.47 | 8.67 |
| CNCN– | 3.44 | 8.64 |
| CNCN– | 3.45 | 8.65 |
| CNCN– | 3.45 | 8.65 |
| CNCN– | 3.42 | 8.62 |
| CNCN– | 3.40 | 8.60 |
| CNCN– | 3.40 | 8.60 |
| CNCN– | 3.41 | 8.61 |
| CNCB– | 4.71 | 4.71 |
| CNCB– | 4.75 | 4.75 |
| CNCB– | 4.84 | 4.84 |
| CNCB– | 5.13 | 5.13 |
| CNCB– | 5.14 | 5.14 |
| CNCB– | 5.15 | 5.15 |
| CNCB– | 5.17 | 5.17 |
The formation energies () of the investigated CXCY defects are presented in Table 2. For CBCB and CNCN defects, is primarily determined by the chemical environment (i.e., the chemical potential of nitrogen). Under their most favorable conditions –N-rich for CBCB and N-poor for CNCN– the formation energies for CBCB configurations are found to be in the narrow range of 3.85-3.88 eV, while for CNCN configurations they range from 2.88 eV to 3.47 eV. Conversely, in their least favorable environments –N-poor for CBCB and N-rich for CNCN– increases substantially, to approximately 9.05-9.08 eV for CBCB and 8.08-8.67 eV for CNCN.
Generally, for these homonuclear pairs (CBCB and CNCN), variations in the specific C–C distance () or interlayer arrangement () have a relatively minor impact on (typically less than 0.05 eV), especially when compared to the large shifts induced by the chemical environment. A prominent exception to this trend is the CNCN–(1,1) configuration. It exhibits the lowest formation energy among all CNCN defects under both N-poor (2.88 eV) and N-rich (8.08 eV) conditions. This enhanced stability, representing a reduction of approximately 0.5-0.6 eV compared to other CNCN configurations at , is consistent with the formation of a stabilizing covalent C–C bond, as discussed in relation to its structure and spin state.
For the heteronuclear CNCB defects, the calculated formation energies (Table 2) range from approximately 4.71 eV to 5.17 eV. These formation energies are identical under N-poor and N-rich conditions because the creation of a CNCB pair involves substituting one B and one N atom with two C atoms; thus, the terms related to and in the formation energy expression effectively cancel each other. The formation energies for the inter-layer CNCB configurations investigated here (all with or , and not forming direct covalent C-C bonds between the pair) are substantially larger than the eV reported for an adjacent CN-CB pair within the same hBN layer () [32]. The significantly lower energy of this intra-layer C2 dimer is likely attributable to the formation of a direct C–C covalent bond, a strong stabilizing effect similar to that observed for the CNCN–(1,1) configuration in our study.
For broader context, other types of carbon defects, such as topological CNCB defects (e.g., Stone-Wales-like structures), can exhibit even higher formation energies than the substitutional CNCB pairs in Table 2, although their can be significantly reduced by mechanical strain [68]. Indeed, strain engineering through biaxial stress has been shown to modulate the formation energies of individual CB or CN defects [69]. While large, uniform strains are not typically expected in bulk hBN, localized stresses, for instance near grain boundaries, could influence defect formation and stability, with carbon defects potentially acting to relieve such stresses [68].
The overall thermal stability and persistence of these C defects in hBN are governed not only by their formation energies but also critically by the kinetic barriers against their migration or transformation. For instance, Babar et al. [68] calculated a high energy barrier (over 6 eV) for the transformation between a Stone-Wales C2 dimer and a standard substitutional C2 dimer, highlighting that C-C bond rearrangements can be significantly hindered. While specific migration and transformation pathways for the CXCY–() configurations in our study have not been computed, similar considerations of strong covalent bond stability suggest that their diffusion and spontaneous restructuring are also likely to be impeded by substantial energy barriers, limiting thermally induced migration under typical experimental conditions.
To further underscore the expected kinetic stability of these C–C pairs, we can estimate the energy cost associated with a specific decomposition mechanism. Consider a CBCB pair, for instance, CBCB–(). One possible route for its partial decomposition involves one of the carbon atoms transitioning from its boron substitutional site (leaving it as VB) into an interstitial position (Ci), while the other carbon atom remains as a CB. The products of such a process would be an isolated CB defect, a carbon interstitial Ci, and a boron vacancy VB. We can estimate the formation energy of this dissociated state, , using the formation energies of these individual neutral defects.
These values render the proposed decomposition pathway highly improbable, even at high temperature (i.e., a barrier over 10 eV). While is about one half of the formation energy of a CB dimer, the interstitial and the vacancy have a much higher , which is expected, as in both defects, strong covalent bonds become dangling bonds. Beyond such decomposition mechanisms, local structural transformations could also occur. For example, an in-plane rotation of a CB-CN bond pair, a process analogous to a step in Stone-Wales defect formation, would constitute another pathway for atomic rearrangement. Again, covalent bonds need to be broken and reformed, and a very high barrier is expected for such processes in hBN [68].
A similar rationale would apply to the decomposition of CNCN pairs, provided the formation energies of the analogous dissociated products (e.g., CN, Ci, VN) result in a similarly high overall energy cost. This analysis, while specific to one decomposition channel, supports the general expectation that transforming these substitutional carbon pairs into defect configurations involving interstitials and vacancies is kinetically strongly hindered. Overall, the robustness of the covalent network in hBN implies that most pathways for significant structural alteration or migration of these carbon defect pairs will involve high activation energies.
The ZPL energy for the defects studied is shown in Fig. 3, as calculated with the SCAN XC functional. In the supplementary material, Fig. S1 shows the values calculated with HSE06 (see the Computational Methods). The donor-acceptor-like dimers have a ZPL energy decreasing with the C–C distance. As the individual defects are charged, their electrostatic interaction is considerable even at long distances. The emission should be in the green-orange range of the visible spectrum. The use of HSE06 is inconsistent for these heteronuclear defects, as we used a different thickness for , effectively changing the dielectric constant.
Defect dimers of the same type, CNCN and CBCB, have a ZPL practically independent of the C–C distance (the smallest distance considered is over 3.5 Å). For CBCB defects, the ZPL is close to eV, and for CNCN it is close to eV, see Fig. 3. To have a reference value, with the HSE06 hybrid functional (Fig. 1 in the supplementary information), the ZPL energies of CBCB and CNCN are 2.1 and 2.4 eV, respectively. Again, the values are almost independent of the C–C distance. It is worth mentioning that, as there is no strong electrostatic interaction, this time the values of the HSE06 functional are not heavily affected by changing the dielectric constant. It is well documented that DFT has an inherent problem in determining accurately the band gap of materials [70]. This is particularly true for defect states [71]. Then, additional information, such as the vibrational spectra, is useful to determine the defect nature against experimental data [40].
The CNCN–(1,1) configuration, which is characterized by the formation of a direct C-C covalent bond, was considered a distinct case. Its optical electronic transition involving this strong C-C bond (e.g., ) should occur at energies larger than the band gap of hBN. For this reason, its ZPL associated with such deep defect states was not computed.
Two types of photoluminescence spectra were found, see Fig. 4A. The one associated with the defect CBCB– represents most of the defects studied here. Figures S2–S4 in the supplementary material show the photoluminescence spectra of all the studied defects emitting within the visible range. These high-energy phonon replicas are typically found in C-based defects in hBN [40, 1]. They correspond to (in-plane) bond stretching involving the C atoms and their closest neighbors, see Fig. 4B. While so far, only defects and defect clusters within a single layer had been considered, the mechanism behind the activation of this phonon mode should apply regardless of the exact location of both C atoms: one electron jumps between defect atoms, or from/to a defect level and the host bands, strengthening or weakening some bonds and causing the in-plane stretching.
@@
However, there are a couple of exceptions, CBCB– (shown in Fig. 4C), and CBCB– with prominent phonon replicas of meV. In these cases, an out-of-plane phonon mode resonates with the transition. In the next section, we will discuss the origin of both types of phonon sidebands found.
@@ IV.1 Triplet or singlet spin ground state
The relation between the spin state and the type of defect pair can be understood by examining the energy and nature of the defect levels within the fundamental band gap of hBN. As established, CB sites introduce donor-like defect states, positioned closer to the conduction band minimum. In contrast, CN sites create acceptor-like defect states, situated nearer to the valence band maximum. This distinction is illustrated schematically in panels A (for a CNCB type) and D (for a CBCB type) of Fig. 5, using the CBCN– and CBCB– configurations as respective examples.
Detailed electronic band structures for many of the defect configurations studied are provided in Figures S5-S7 of the Supplementary Material. For cases where singlet and triplet states of a defect pair were found to be nearly degenerate in energy, only the band structure for the triplet state is presented, as the singlet counterpart yields almost identical results. Generally, the donor/acceptor picture for CB/CN holds as long as the carbon defects maintain their sp2 hybridization within the hBN plane. A notable exception is the CNCN– configuration (see Fig. S7 in the Supplementary Material), which undergoes significant structural rearrangement to form a C-C bond with sp3-like character. For this defect, a single, non-degenerate defect-related band appears within the hBN band gap, with other associated levels located deep within the valence band.
For heteronuclear CNCB defect pairs, the inherent donor (CB) and acceptor (CN) nature of the constituent defects leads to charge transfer from CB to CN, resulting in C and C ionic states. This is depicted in Fig. 5A, where the occupied defect levels (associated with C) are shown lower in the gap, and empty levels (associated with C) are higher. The corresponding wavefunctions (Fig. 5B,C) illustrate the spatial localization of these occupied and empty states primarily on the CN and CB sites, respectively. This charge transfer results in a doubly occupied acceptor state localized on the CN site (forming C) and an emptied donor site (forming C). As these resulting ionic centers (C and C) are effectively non-magnetic due to their spin-paired electronic configurations, the CNCB defect pair consequently exhibits a singlet ground state.
In contrast, for homonuclear defect pairs, such as CBCB (or CNCN), both components are of the same type (e.g., both donors for CBCB). For these homonuclear pairs, the identical nature of the two defects means their respective unperturbed defect states are intrinsically degenerate. When these defects interact at sufficiently short distances, this degeneracy is lifted, and new combined states are formed which can be described as linear combinations of the original defect states, as illustrated for CBCB in Fig. 5D-F. For these systems, each C atom contributes an electron to the defect-related manifold. The significant overlap between the wavefunctions of these occupied levels gives rise to an exchange interaction. This interaction typically stabilizes the triplet state (parallel spins), where the spatial wavefunction is antisymmetric, relative to the singlet state (antiparallel spins), which has a symmetric spatial wavefunction, thus making the triplet state the ground state at shorter distances. As the separation between the identical defects increases, the wavefunction overlap and the consequent exchange interaction diminish. In this limit, the energy difference between the singlet and triplet states vanishes, leading to their degeneracy as the system effectively becomes two weakly interacting centers.